Calculate the tax disadvantage to organizing a U.S. business as a corporation versus a partnership, given the following assumptions. All earnings will be paid out as dividends, and operating income before taxes will be $1,869,564. The effective corporate tax rate is 21%, and the tax rate on corporate dividends is 15%. The average personal tax rate for partners in the business is 31%.
Answer #1

Net Income After Tax in the hands of partners:

For Corporate:

(EBIT - Corporate Tax)/(1+Corporate Dividends Tax Rate)

= (1869564 - 21%)/(1+0.15)

= 1476955.56/1.15 = 1284309.1826

For Partnership:

EBIT - Tax

1869564 - 31% = 1289999.16

Therefore, Tax Disadvantage for Corporate = Income as per Partnership - Income as per Corporate

= 1289999.16 - 1284309.1826 = 5689.5774

(If this was helpful then please rate positively. Thank You:)